# A Unified Pulse-Shaped OFDM Framework for Chirp-Domain Waveforms: Continuous-Time Modeling and Practical I/O Analysis

### 15 Mar 2026 v1

@@ Appendix B Derivation of the Effective Channel Tap $h_{k^{\prime},l}^{\mathrm{MF}}$
The channel tap $h_{k^{\prime},l}^{\mathrm{MF}}$ can be interpreted as the channel observed when a discrete sequence with symbol interval $T_{s}$, pulse-shaped by $a(t)$ with Nyquist interval $T_{s}$, passes through the DD channel and is subsequently processed by a matched filter.
Consider the transmitted waveform
$x(t)=\sum_{k}x[k]a\left(t-kT_{s}\right).$ (60)
Substituting (60) into (41) and ignoring the noise term yields
$y(t)$ $=\sum_{p=1}^{P}h_{p}e^{j2\pi\nu_{p}(t-\tau_{p})}\sum_{k}x[k]a(t-\tau_{p}-kT_{s}),$
$=\sum_{p=1}^{P}\sum_{k}h_{p}x[k]e^{j2\pi\nu_{p}(t-\tau_{p})}a(t-\tau_{p}-kT_{s}).$
The output of the matched filter is given by
@@
$\quad\times\int_{-\infty}^{\infty}a(\xi)a^{*}(\xi-(t-\tau_{p}-kT_{s}))e^{j2\pi\nu_{p}\xi}d\xi.$
Letting $\tau=t-\tau_{p}-kT_{s}$, we obtain
$y^{\mathrm{MF}}(t)$ $=\sum_{p=1}^{P}\sum_{k}h_{p}x[k]e^{j2\pi\nu_{p}kT_{s}}\mathcal{A}_{a}(t-\tau_{p}-kT_{s},\nu_{p}),$
where $\mathcal{A}_{a}(\tau,\nu)$ is the ambiguity function of $a(t)$.
Next, sampling at $t=\tau_{1}+k^{\prime}T_{s}$ and defining the delay tap index $l\triangleq k^{\prime}-k$, we obtain
@@
$=\sum_{l}\sum_{p=1}^{P}h_{p}e^{j2\pi\nu_{p}(k^{\prime}-l)T_{s}}$
$\quad\times\mathcal{A}_{a}(lT_{s}+\tau_{1}-\tau_{p},\nu_{p})x[k^{\prime}-l].$
@@
$h_{k^{\prime},l}^{\mathrm{MF}}\triangleq\sum_{p=1}^{P}h_{p}e^{j2\pi\nu_{p}(k^{\prime}-l)T_{s}}\mathcal{A}_{a}(lT_{s}+\tau_{1}-\tau_{p},\nu_{p}).$
@@ Appendix C Derivation of Channel Matrix $\mathbf{H}^{\rm MF}_{u}$
Since $A_{a}(\cdot)$ has a finite support, we may limit the delay index $l$ as $l=0,\cdots,L-1$ and rewrite (49) as
@@
$\hat{\mathbf{H}}^{\mathrm{MF}}$ $=\hat{\mathbf{H}}^{\mathrm{MF}}\mathbf{x}^{\prime}+\mathbf{z}^{\mathrm{MF}},$ (62)
@@
$[\hat{\mathbf{H}}^{\mathrm{MF}}]_{k^{\prime},n}$ $=\begin{cases}h_{k^{\prime},k^{\prime}-n}^{\mathrm{MF}},&0\leq k^{\prime}-n\leq L-1,\\ 0,&\text{otherwise},\end{cases}$ (63)
for $k^{\prime}=0,\cdots,N-1$ and $n=-L+1,\cdots,N-1$, and the negative-index entries in $\mathbf{x}^{\prime}$ corresponds to the CPP.
@@
$h_{k^{\prime},l}^{\mathrm{MF}\prime}$ $\triangleq h_{k^{\prime},l}^{\mathrm{MF}}e^{-j2\pi c_{1}\left(N^{2}+2N(k^{\prime}-l)\right)}.$ (64)
@@
$+\sum_{l=k^{\prime}+1}^{L-1}h_{k^{\prime},l}^{\mathrm{MF}\prime}x_{\rm AF}[N+k^{\prime}-l]+z^{\mathrm{MF}}[k^{\prime}].$
Hence, let $\mathbf{x}=[x_{\rm AF}[0],x_{\rm AF}[1],\cdots,x_{\rm AF}[N-1]]^{\mathsf{T}}$, we obtain the CPP-folded square-matrix representation
$\mathbf{y}^{\mathrm{MF}}$ $=\mathbf{H}^{\rm MF}\mathbf{x}+\mathbf{z}^{\mathrm{MF}},$ (65)
@@
$[\mathbf{H}^{\rm MF}]_{k^{\prime},N+k^{\prime}-l}$ $=h_{k^{\prime},l}^{\mathrm{MF}\prime},\qquad l=k^{\prime}+1,\cdots,L-1.$ (66)
@@
where the $(q,k)$-th entry of $\mathbf{H}_{u}^{\mathrm{MF}}$ is shown on the top of the next page.
$[\mathbf{H}_{u}^{\mathrm{MF}}]_{q,k}$ $=\sum_{k^{\prime}=0}^{N-1}\sum_{n=0}^{N-1}\phi^{*}[k^{\prime},q]\,[\mathbf{H}^{\mathrm{MF}}]_{k^{\prime},n}\,\phi[n,k]$
$=\sum_{k^{\prime}=0}^{N-1}\phi^{*}[k^{\prime},q]\left(\sum_{l=0}^{k^{\prime}}h_{k^{\prime},l}^{\mathrm{MF}}\phi[k^{\prime}-l,k]+\sum_{l=k^{\prime}+1}^{L-1}h_{k^{\prime},l}^{\mathrm{MF}\prime}\phi[N+k^{\prime}-l,k]\right)$
$=\frac{1}{N}e^{j2\pi c_{2}(k^{2}-q^{2})}\sum_{k^{\prime}=0}^{N-1}\sum_{l=0}^{L-1}h_{k^{\prime},l}^{\mathrm{MF}}e^{j2\pi c_{1}\left((k^{\prime}-l)^{2}-k^{\prime 2}\right)}e^{j2\pi\left(\frac{(k^{\prime}-l)k-k^{\prime}q}{N}\right)}.$ (68)

### 31 Mar 2026 v2

@@ Appendix B Derivation of the Effective Channel Tap $h_{k^{\prime},l}^{\mathrm{MF}}$
The channel tap $h_{k^{\prime},l}^{\mathrm{MF}}$ can be interpreted as the channel observed when a discrete sequence with sample interval $T_{s}$, pulse-shaped by $a(t)$, passes through the DD channel and is subsequently processed by a matched filter.
Let us consider the transmitted waveform $x(t)=\sum_{k}x[k]a\left(t-kT_{s}\right)$, and substitute it into (41) to yield
$y(t)$ $=\sum_{p=1}^{P}\sum_{k}h_{p}x[k]e^{j2\pi\nu_{p}(t-\tau_{p})}a(t-\tau_{p}-kT_{s}),$
where the noise term is ignored. The output of the matched filter is given by
@@
$\quad\times\int_{-\infty}^{\infty}a(\xi)a^{*}(\xi-(t-\tau_{p}-kT_{s}))e^{j2\pi\nu_{p}\xi}d\xi,$
$=\sum_{p=1}^{P}\sum_{k}h_{p}x[k]e^{j2\pi\nu_{p}(t-\tau_{p})}\mathcal{A}_{a,a}(\hat{\tau},-\nu),$
where $\hat{\tau}=t-\tau_{p}-kT_{s}$. Then, sampling at $t=\tau_{1}+k^{\prime}T_{s}$ and defining the delay tap index $l\triangleq k^{\prime}-k$, we obtain
@@
$=\sum_{l}\sum_{p=1}^{P}h_{p}e^{j2\pi\nu_{p}(k^{\prime}T_{s}-\tilde{\tau}_{p})}$
$\quad\times\mathcal{A}_{a,a}(lT_{s}-\tilde{\tau}_{p},-\nu_{p})x[k^{\prime}-l].$
@@
$h_{k^{\prime},l}^{\mathrm{MF}}\triangleq\sum_{p=1}^{P}h_{p}e^{j2\pi\nu_{p}(k^{\prime}T_{s}-\tilde{\tau}_{p})}\mathcal{A}_{a,a}(lT_{s}-\tilde{\tau}_{p},-\nu_{p}).$
@@ Appendix C Derivation of Channel Matrix $\mathbf{H}^{\rm MF}_{u}$
Since $A_{a,a}(\cdot)$ has a finite support, we may limit the delay index $l$ as $l=0,\cdots,L-1$ and rewrite (52) as
@@
$\mathbf{y}^{\mathrm{MF}}$ $=\hat{\mathbf{H}}^{\mathrm{MF}}\mathbf{x}^{\prime}+\mathbf{z}^{\mathrm{MF}},$ (63)
@@
$[\hat{\mathbf{H}}^{\mathrm{MF}}]_{k^{\prime},\dot{n}}$ $=\begin{cases}h_{k^{\prime},k^{\prime}-\dot{n}}^{\mathrm{MF}},&0\leq k^{\prime}-\dot{n}\leq L-1,\\ 0,&\text{otherwise},\end{cases}$ (64)
for $k^{\prime}=0,\cdots,N-1$ and $\dot{n}=-L+1,\cdots,N-1$, and the negative-index entries in $\mathbf{x}^{\prime}$ corresponds to the CPP, which is given by
$x_{\rm AF}[k]=x_{\rm AF}[N+k]e^{-j2\pi c_{1}(N^{2}+2Nk)}$ (65)
for $k=-L_{\rm CPP},\dots,-1$ and $L_{\rm CPP}\geq L$.
@@
$\acute{h}_{k^{\prime},l}^{\mathrm{MF}}$ $\triangleq h_{k^{\prime},l}^{\mathrm{MF}}e^{-j2\pi c_{1}\left(N^{2}+2N(k^{\prime}-l)\right)}.$ (66)
@@
$+\sum_{l=k^{\prime}+1}^{L-1}\acute{h}_{k^{\prime},l}^{\mathrm{MF}}x_{\rm AF}[N+k^{\prime}-l]+z^{\mathrm{MF}}[k^{\prime}].$
Hence, let $\mathbf{x}_{\rm AF}=[x_{\rm AF}[0],x_{\rm AF}[1],\cdots,x_{\rm AF}[N-1]]^{\mathsf{T}}$, we obtain the CPP-folded square-matrix representation
$\mathbf{y}^{\mathrm{MF}}$ $=\mathbf{H}^{\rm MF}\mathbf{x}_{\rm AF}+\mathbf{z}^{\mathrm{MF}},$ (67)
@@
$[\mathbf{H}^{\rm MF}]_{k^{\prime},N+k^{\prime}-l}$ $=\acute{h}_{k^{\prime},l}^{\mathrm{MF}},\qquad l=k^{\prime}+1,\cdots,L-1.$ (68)
@@
where the $(n^{\prime},n)$-th entry of $\mathbf{H}_{u}^{\mathrm{MF}}$ is shown at the top of the next page.
$[\mathbf{H}_{u}^{\mathrm{MF}}]_{n^{\prime},n}$ $=\sum_{k^{\prime}=0}^{N-1}\sum_{\dot{n}=0}^{N-1}[\mathbf{A}]_{n^{\prime},k^{\prime}}[\mathbf{H}^{\mathrm{MF}}]_{k^{\prime},\dot{n}}[\mathbf{A}^{\mathsf{H}}]_{\dot{n},n}$
$=\sum_{k^{\prime}=0}^{N-1}[\mathbf{A}]_{n^{\prime},k^{\prime}}\left(\sum_{l=0}^{k^{\prime}}h_{k^{\prime},l}^{\mathrm{MF}}[\mathbf{A}^{\mathsf{H}}]_{k^{\prime}-l,n}+\sum_{l=k^{\prime}+1}^{L-1}\acute{h}_{k^{\prime},l}^{\mathrm{MF}}[\mathbf{A}^{\mathsf{H}}]_{N+k^{\prime}-l,n}\right)$
$=\frac{1}{N}e^{j2\pi c_{2}(n^{2}-{n^{\prime}}^{2})}\sum_{k^{\prime}=0}^{N-1}\sum_{l=0}^{L-1}h_{k^{\prime},l}^{\mathrm{MF}}e^{j2\pi c_{1}\left((k^{\prime}-l)^{2}-{k^{\prime}}^{2}\right)}e^{j2\pi\left(\frac{(k^{\prime}-l)n-k^{\prime}n^{\prime}}{N}\right)}$ (70)